OpenStudy (chrisneedshelp):

3(2x - 3) + 4(2 - x) = 0 1/2 17/10 -17/2

OpenStudy (whpalmer4):

\[3(2x-3) + 4(2-x) = 0\]Can you use the distributive property to remove the parentheses?

OpenStudy (anonymous):

3*2x=(3*2)x=6x 3*3=(3*-3)=-9 6x-9 4*2=(4*2)= 8 4*-x=(4*-1)x=-4x 8-4x 6x-9+4x=-4x+8+4x 10x-9=8 10x-9+9=8+9 10x=17 10x/10=17/10 x=1.7

OpenStudy (whpalmer4):

Sorry, but that's incorrect...

OpenStudy (chrisneedshelp):

lol i was eating

OpenStudy (chrisneedshelp):

those are fractions in the end

OpenStudy (chrisneedshelp):

i dont think you can use the distributive property for the parenthesis

OpenStudy (anonymous):

the end is dividing both sides of the equation by 10

OpenStudy (whpalmer4):

you most certainly can use the distributive property! a(b+c) = a*b + a*c a(b-c) = a*b - a*c

OpenStudy (whpalmer4):

@ken_boy05 your mistake is in this line \[6x-9+4x=-4x+8+4x\]

OpenStudy (whpalmer4):

@ChrisNeedsHelp stick to the problem we're working on, eh?

OpenStudy (anonymous):

you need to group your variables 6x and -4x to do that you add 4x to both sides 6x-9+4x=-4x+8+4x

OpenStudy (whpalmer4):

@ken_boy05 if you still think you are getting the right answer, plug it into the original equation. does it fit?

OpenStudy (anonymous):

so whats your answer@whpalmer4?

OpenStudy (whpalmer4):

I'm waiting for the original poster to do some work, I'm not going to give him the answer.

OpenStudy (anonymous):

lol

OpenStudy (triciaal):

@ken_boy05 he is right you made a mistake

OpenStudy (anonymous):

Well obviously he said he is eating so he doesn't really care that much lol

OpenStudy (triciaal):

6x -9 + 8 - 4x = 0

OpenStudy (anonymous):

Well I grew up in England and moved here from Canada ...the British way sometimes is different ...I could be wrong but that's how I learnt

OpenStudy (anonymous):

tricia that part yes i did not include it was b4 6x-9+4x=-4x+8+4x

OpenStudy (anonymous):

@6x -9 + 8 - 4x = 0

OpenStudy (triciaal):

10x-9+9=8+9 this is where you made the mistake

OpenStudy (triciaal):

it was -9 on the right

OpenStudy (triciaal):

there is nothing wrong with your method!!!

OpenStudy (whpalmer4):

actually there's a mistake in the line I pointed out — wrong sign on the 4x

OpenStudy (whpalmer4):

on the left hand side

OpenStudy (whpalmer4):

4(2-x) = 8 - 4x

OpenStudy (anonymous):

American way is not the only way ...your method i can already tell might end up in x=0.5 and its correct as well

OpenStudy (anonymous):

Math is logic

OpenStudy (anonymous):

Can you please help Chris? Thanks

OpenStudy (whpalmer4):

Chris needs to pull on the big boy pants and do some work. I told him the first step, and will help him further when he does it. There's no American way versus British way or Canadian way here, don't be ridiculous. Algebraic manipulation is the same no matter where you are.

OpenStudy (whpalmer4):

The issue was that you changed a sign when you shouldn't have. An honest mistake, happens all the time.

OpenStudy (triciaal):

agree Algebraic manipulation is the same no matter where you are.

OpenStudy (anonymous):

so your argument is the manipulation is only one way ??smh ...The ancient Arabs and Egyptians that started math write and work from left to right ...plsssss...show him ur method and we already know its x=0.5

OpenStudy (whpalmer4):

Good reason for checking your work before posting it :-)

OpenStudy (whpalmer4):

produce a living ancient Arab or Egyptian, and I'll be happy to discuss it with them.

OpenStudy (whpalmer4):

the underlying mathematics do not change just because you use Roman numerals or Babylonian notation.

OpenStudy (anonymous):

The ancient manuscripts are still alive and healthy ...you haven't seen one before

OpenStudy (whpalmer4):

the distributive principle does not depend on the year or the lat/long

OpenStudy (anonymous):

Just help Chris pls lol Thanks

OpenStudy (whpalmer4):

Ken, you made a mistake. Don't try to blame dead people for it.

OpenStudy (anonymous):

well you that is alive don't want to help chris

OpenStudy (whpalmer4):

Chris can only be helped if he takes part.

OpenStudy (anonymous):

3(2x-3) + 4(2-x) = 0 3*2x - 3*3 + 4*2 + 4(-x) = 0 6x - 9 + 8 - 4x = 0 6x - 1 - 4x = 0 2x - 1 = 0 2x = 1 x = 1/2 or 0.5

OpenStudy (triciaal):

enough by the way,I learnt the "British" way and a mistake is just that

OpenStudy (anonymous):

Chris have a great night ...that's your American answer

OpenStudy (anonymous):

Still cant find clue to this maybe you can help Find an equation in point -slope form for the line having the specified slope and containing the point indicated m=4, ( 3,8 )

OpenStudy (anonymous):

@whpalmer4 and triciaal

OpenStudy (whpalmer4):

@ken_boy05 please don't put new questions in someone else's question. post it as a new question and I'll be there with the answer for you.

OpenStudy (anonymous):

i put it there before already @whpalmer4